a triangle with a perimeter of 40 meters is 2 meters longer than its wide. Deterrmine the demensions.
Answer by macston(5194)

W=width; L=length=W+2m; P=perimeter=2(L+W)=40m
P=2(L+W)
40m=2(L+W) Divide each side by 2.
20m=L+W Substitute for L.
20m=W+2m+W Subtract 2m from each side.
18m=2W Divide each side by 2.
9m=W ANSWER 1: The width is 9 meters.
L=W+2m=9m+2m=11m ANSWER 2: The length is 11 meters
CHECK:
P=2(L+W)
40m=2(11m+9m)
40m=2(20m)
40m=40m